A cylinder of mass m and radius R has a moment of inertia of 12mr2. The cylinder is released from rest at a height h on an inclined plane, and rolls down the plane without slipping. What is the velocity of the cylinder when it reaches the bottom of the incline?





1
Expert's answer
2022-02-01T09:56:43-0500

Given:

"I=1\/2mR^2"


The law of conservation of the energy says

"E_i=E_f"

"E_{ki}+E_{pi}=E_{kf}+E_{pf}"

"mgh=\frac{mv_f^2}{2}+\frac{I\omega_f^2}{2}""\omega_f=v_f/R"

Hence

"mgh=\frac{mv_f^2}{2}+\frac{mv_f^2}{4}=\frac{3mv_f^2}{4}"

Finally, the velocity of the cylinder when it reaches the bottom of the incline

"v_f=\sqrt{4/3gh}"